OpenStudy (anonymous):

(3x^2y^5z^-2 / 5xz^5)^-3 PLEASE HELP! I understand how to do everything except the negative exponent outside the parentheses.

OpenStudy (anonymous):

multiply each exponent by \(-3\) first

OpenStudy (anonymous):

a^-3 = (1/a)^3 so invert and cube (a^n)^m = a^nm powers multiply